These are the intercepts. The intervals between the intercepts are called the axes of this ellipse; the larger (resp. smaller) is called the major axis (resp. minor axis); thus the larger of a and b is the semi-major axis while the smaller is the semi-minor axis.

Index

axis, 1, ellipse1 axis , -major 1, -minor 1 intercepts, 1